A 2 nm sized cube containing an unknown distribution of charges is sitting inside a vacuum. The magnitude of the repulsive force acting on an electron 1 cm away from the cube is measured to be 1.7x10-22 N. (Hint: 1nm = 10-9 m)

a. Estimate the amount of charges inside the box.

b. About how many electrons does that charge correspond to?
Answer #1

a) F = kqe/r^2 ======> q = F*r^2/ke

     q = (1.7*10^-22*10^-4)/(9*10^9*1.6*10^-19) = 1.18*10^-17 C

b) q = ne ========> n = q/e = 1.18*10^-17/1.6*10^-19 = 73.75 ~ 74 electrons